OpenStudy (anonymous):

What is the midpoint of the line segment with endpoints (5, -2) and (1, -6)?

OpenStudy (anonymous):

Have you tried drawing it? I guess, that would really help..

OpenStudy (anonymous):

Midpoint formula is \[Midpoint = (\frac{x1+x2}{2}, \frac{y1+y2}{2})\] \[=(\frac{5+1}{2}, \frac{-2+-6}{2}) \] \[= (\frac{6}{2}, \frac{-8}{2})\] \[=(3,-4)\]

OpenStudy (anonymous):

Would be much better, if @saheiiiim actually understood, where this "formula" of yours comes from, I think - so please don't just post answers @JayDS. It doesn't help much and it also violates the OpenStudy code of conduct. Thanks :)